What is the Y intercept of the line given by this equation 3x-y=6
Answer by hovuquocan1997(83)

If you want to find y-intercept, you have to set the x in the equation to 0
3x - y = 6
3*0 - y = 6
-y = 6
y = -6
So your y intercept is (0,-6)
TA-DAH
:D